# Edge open packing on subclasses of chordal graphs

### v1

@@ 1 Introduction
In this paper, we consider $G$ as a finite simple graph with vertex set $V(G)$ and edge set $E(G)$. Any terms and notation not explicitly defined here follow the definitions in [16]. A graph $H=(V^{\prime},E^{\prime})$ is called a subgraph of a graph $G=(V,E)$ if $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$. Given a vertex subset $S\subseteq V$, the induced subgraph $G[S]$ is the subgraph with vertex set $S$ and edge set consisting precisely of those edges in $E$ whose endpoints are both in $S$. Similarly, for an edge subset $B\subseteq E(G)$, the induced subgraph $G[B]$ has a vertex set comprising all endpoints of edges in $B$ and an edge set exactly $B$ itself. We use the notation $G\langle D\rangle$ to explicitly denote the edge-induced subgraph on a set of edges $D\subseteq E(G)$. The open neighbourhood of a vertex $x\in V(G)$, denoted $N_{G}(x)$, is the set of vertices adjacent to $x$. The closed neighbourhood of $x$, denoted $N_{G}[x]$, is defined by $N_{G}[x]=N_{G}(x)\cup{x}$. The minimum and maximum degrees of $G$ are denoted by $\delta(G)$ and $\Delta(G)$, respectively. For a connected graph $G=(V,E)$, the distance between two vertices $u,v\in V$, denoted by $d_{G}(u,v)$, is the length (number of edges) of a shortest path connecting them. The diameter of $G$, written as $\operatorname{diam}(G)$, is the maximum distance between any two vertices in $G$, formally defined as $\operatorname{diam}(G)=\displaystyle\max_{u,v\in V}d_{G}(u,v)$. A graph is called bipartite if its vertex set can be divided into two disjoint independent sets. A star is the complete bipartite graph $K_{1,t}$ for $t\geq 1$.
@@
An induced matching $M$ of a graph $G$ is a subset $M\subseteq E(G)$ such that $M$ constitutes a matching and no two edges in $M$ are connected by any edge of $G$. The induced matching number $G$ represents the maximum size of an induced matching within $G$. The problem of finding a maximum induced matching was first presented by Stockmeyer and Vazirani as the “risk-free marriage problem” [15] and has since been thoroughly examined (see, for instance, [4, 7, 11]).
@@
In 2022, Chelladurai et al. [6] introduced the concept of an edge open packing (EOP) in graphs. A subset $D\subseteq E(G)$ is termed an edge open packing set (EOP set) if, for every pair of distinct edges $e_{1},e_{2}\in D$, no edge in $E(G)\setminus{e_{1},e_{2}}$ connects endpoints of $e_{1}$ with endpoints of $e_{2}$. Equivalently, for any two edges $e_{1},e_{2}\in D$, there is no third edge $e_{3}\in E(G)\setminus{e_{1},e_{2}}$ forming either a path $P_{4}$ (with four vertices) or a cycle $C_{3}$ (with three vertices) in the subgraph $G\langle e_{1},e_{2},e_{3}\rangle$. The edge open packing number, denoted by $\rho^{o}_{e}(G)$, is defined as the maximum size of an EOP set in $G$. An EOP set with cardinality equal to $\rho^{o}_{e}(G)$ is called a $\rho^{o}_{e}$-set. It is noteworthy that if $D$ is an EOP set (resp. an induced matching), the induced subgraph $G[D]$ is a disjoint union of induced stars (resp. $K_{1,1}$-stars). Thus, the induced matching problem can be viewed as a special case of the EOP problem, which relaxes the stricter conditions of a matching or induced matching.
@@
An injective $k$-edge coloring of a graph $G$ is an assignment of colors (integers) from the set $\{1,\ldots,k\}$ to the edges of $G$ such that any two edges $e_{1},e_{2}\in E(G)$ that share a common adjacent edge, meaning an edge $e$ exists with one endpoint common to $e_{1}$ and the other endpoint common to $e_{2}$, are assigned distinct colors. The smallest integer $k$ for which this coloring exists is known as the injective chromatic index of $G$, denoted by $\chi^{\prime}_{i}(G)$. This notion was introduced by Cardoso et al. [5] and subsequently explored in greater depth in [8, 12]. It is noteworthy that each color class generated by an injective edge coloring constitutes an edge open packing (EOP) set. Consequently, determining $\chi^{\prime}_{i}(G)$ is effectively the problem of partitioning the edge set $E(G)$ into EOP sets. Practical applications of injective edge colorings and edge open packings have been discussed in the literature, specifically in [5] and [6].
@@
In [6], Chelladurai et al. studied edge open packing in graphs. They showed some bounds in terms of a graph’s diameter, size, minimum degree, clique number, and girth. Also, they characterized graphs $G$ with $\rho_{e}^{o}(G)\in\{1,2,m-2,m-1,m\}$, where $m$ is the number of edges of $G$. Further, Brešar and Samadi [3] study the EOP set problem from an algorithmic point of view. They proved that the decision problem associated with Maximum Edge Open Packing Problem is NP-complete for three special families of graphs, namely, graphs with universal vertices, Eulerian bipartite graphs, and planar graphs with maximum degree $4$. Moreover, they designed a linear-time algorithm in trees to solve the Maximum Edge Open Packing Problem. Also, they provided a structural characterization of all graphs $G$ attaining the upper bound $\rho_{e}^{o}(G)\leq|E(G)|/\delta(G)$. They analyzed the effect of edge removal on the EOP number of a graph $G$, obtaining a lower and an upper bound for $\rho_{e}^{o}(G-e)$, where $e$ is an edge of $G$. In [2], Brešar et al. showed some relationship between the induced matching number and the EOP number in trees and graph products. Later, Pandey and Santra [14] further characterize this parameter for any $t\geq 3$. They showed the necessary and sufficient condition for $\rho_{e}^{o}=t$, for $t\geq 3$, and explicitly defined graph families for $\rho_{e}^{o}=m-3$, where $m$ is the number of edges of $G$. In[3], Brešar and Samadi pose the following open question of the edge open packing problem in chordal graphs.
@@
An edge between two non-consecutive vertices of a cycle is called a chord. If every cycle in $G$ of length at least four has a chord, then $G$ is called a chordal graph. A vertex $v\in V$ is called a simplicial vertex of $G$ if $N_{G}[v]$ induces a clique in $G$. A perfect elimination ordering (PEO) of $G$ is an ordering of the vertices, say $\sigma=(v_{1},v_{2},\ldots,v_{n})$, such that $v_{i}$ is a simplicial vertex of $G_{i}=G[\{v_{i},v_{i+1},\ldots,v_{n}\}]$ for all $1\leq i\leq n$. The existence of PEO can characterize chordal graphs; that is, a graph $G$ is chordal if and only if $G$ has a PEO [9].
@@
For a chordal graph $G$, a perfect elimination ordering (PEO) $\alpha=(v_{1},v_{2},\ldots,v_{n})$ is called a bi-compatible elimination ordering (BCO) if the reverse ordering $\alpha^{-1}=(v_{n},v_{n-1},\ldots,v_{1})$ is also a PEO of $G$. A graph $G$ admits a BCO if and only if it is a proper interval graph [10]. The following observation concerns the structure of BCOs in proper interval graphs.
Observation 2 ([13]).
@@ 2 EOP set in proper interval graphs
Input: A connected proper interval graps $G=(V,E)$ with $n$ vertices.
@@
2Calculate a BCO $\sigma=\{v_{1},v_{2},\ldots,v_{n}\}$ of $G$;
3
4Calculate $v_{k}=l(v_{1})$;
5
6for $j=2$ to $k$ do
7
 $\rho_{e}^{o}\left(\mathcal{H}_{j}\right)$= MaxEOP_PIG($\mathcal{H}_{j}$) ;
 // where $\mathcal{H}_{j}=G[v_{l(v_{j})+1},v_{l(v_{j})+2},\ldots,v_{n}]$
8
9 if $l(v_{j})=k$ then
@@ 11    $j\leftarrow j+1$;
14 else
15
16    for $k<p\leq l(v_{j})$ do
@@
       $\rho_{e}^{o}\left(\mathcal{G}^{p}_{j}\right)$=MaxEOP_PIG($\mathcal{G}^{p}_{j}$) ;
       // where $\mathcal{G}^{p}_{j}=G[v_{l(v_{p})+1},v_{l(v_{p})+2},\ldots,v_{n}]$
18
19
20    $j\leftarrow j+1$;
21
@@ 20    $j\leftarrow j+1$;
23
[…]
25
[…]
27
[…]
29
$\rho_{2}\left(G\right)=\rho_{e}^{o}\left(\mathcal{G}^{\prime}\right)$ ;
// where $\mathcal{G}^{\prime}=G[v_{2},v_{3},\ldots,v_{n}]$
30
[…]
32
33return $\rho_{e}^{o}\left(G\right)$;
34
@@ 3 EOP set in block graphs
In this section, we design a polynomial-time algorithm for finding the edge open packing number of a given block graph. We start with some pertinent definitions related to block graphs. Let $G$ be a block graph with $l$ blocks, $B_{1},B_{2},\ldots,B_{l}$, and $k$ cut vertices, $v_{1},v_{2},\ldots,v_{k}$. The cut-tree of $G$, denoted by $T_{G}=(V^{\prime},E^{\prime})$, is defined as $V^{\prime}=\{B_{1},B_{2},\ldots,B_{l},v_{1},v_{2},\ldots,v_{k}\}$ and $E^{\prime}=\{B_{i}v_{j}|v_{j}\in B_{i},1\leq i\leq l,1\leq j\leq k\}$. An example of a block graph and its cut-tree is illustrated in Figure 1. A vertex of $T_{G}$ corresponding to a block of $G$ is called a block vertex. The cut-tree $T_{G}$ of block graph $G$ can be made a rooted tree by selecting any block or cut vertex as the root. The set of all blocks of a given block graph $G=(V,E)$ can be found in $O(|V|+|E|)$ time by performing a depth-first search (DFS). For each cut vertex $v$ of $G$, the set of all blocks containing $v$ can also be found in $O(|V|+|E|)$ time. Hence the rooted cut-tree of a block graph can be constructed in $O(|V|+|E|)$ time [1].
@@ 5 Conclusion
In this paper, we have partially solved the open question posed by Brešar and Samadi about edge open packing sets in chordal graphs. First, we have designed a polynomial-time algorithm to solve the maximum edge open packing set in a proper interval graph. Furthermore, we have presented a polynomial-time algorithm to solve the maximum edge open packing set in block graphs. Finally, we have shown that the problem of edge open packing number can be solved in linear time in split graphs. It would be fascinating to investigate the other open problems mentioned in [6] and [3].

### v2

@@ 1 Introduction
In this paper, we consider $G$ as a finite simple graph with vertex set $V(G)$ and edge set $E(G)$. Any terms and notation not explicitly defined here follow the definitions in [west2001introduction]. A graph $H=(V^{\prime},E^{\prime})$ is called a subgraph of a graph $G=(V,E)$ if $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$. Given a vertex subset $S\subseteq V$, the induced subgraph $G[S]$ is the subgraph with vertex set $S$ and edge set consisting precisely of those edges in $E$ whose endpoints are both in $S$. Similarly, for an edge subset $B\subseteq E(G)$, $G[B]$ denotes the subgraph of $G$ induced by the endpoints of edges in $B$. Let $D\subseteq E(G)$. An edge-induced subgraph on $D$ of $G$ is defined as the subgraph of $G$ whose edges are the edges from $D$ and the vertices are the endpoints of edges of $D$. This subgraph is denoted by $G\langle D\rangle$. The open neighbourhood of a vertex $x\in V(G)$, denoted $N_{G}(x)$, is the set of vertices adjacent to $x$. The closed neighbourhood of $x$, denoted $N_{G}[x]$, is defined by $N_{G}[x]=N_{G}(x)\cup{x}$. The minimum and maximum degrees of $G$ are denoted by $\delta(G)$ and $\Delta(G)$, respectively. For a connected graph $G=(V,E)$, the distance between two vertices $u,v\in V$, denoted by $d_{G}(u,v)$, is the length (number of edges) of a shortest path connecting them. The diameter of $G$, written as $\operatorname{diam}(G)$, is the maximum distance between any two vertices in $G$, formally defined as $\operatorname{diam}(G)=\displaystyle\max_{u,v\in V}d_{G}(u,v)$. A graph is called bipartite if its vertex set can be divided into two disjoint independent sets. A star is the complete bipartite graph $K_{1,t}$ for $t\geq 1$.
@@
An induced matching $M$ of a graph $G$ is a subset $M\subseteq E(G)$ such that $M$ constitutes a matching and no two edges in $M$ are connected by any edge of $G$. The induced matching number $G$ represents the maximum size of an induced matching within $G$. The problem of finding a maximum induced matching was first presented by Stockmeyer and Vazirani as the “risk-free marriage problem” [stockmeyer1982np] and has since been thoroughly examined (see, for instance, [cameron1989induced, dabrowski2013new, lozin2002maximum]).
@@
In 2022, Chelladurai et al. [chelladurai2022edge] introduced the concept of an edge open packing (EOP) in graphs. A subset $D\subseteq E(G)$ is termed an edge open packing set (EOP set) if, for every pair of distinct edges $e_{1},e_{2}\in D$, no edge in $E(G)\setminus{e_{1},e_{2}}$ connects endpoints of $e_{1}$ with endpoints of $e_{2}$. Equivalently, for any two edges $e_{1},e_{2}\in D$, there is no third edge $e_{3}\in E(G)\setminus{e_{1},e_{2}}$ forming either a path $P_{4}$ (with four vertices) or a cycle $C_{3}$ (with three vertices) in the subgraph $G\langle e_{1},e_{2},e_{3}\rangle$. The edge open packing number, denoted by $\rho^{o}_{e}(G)$, is defined as the maximum size of an EOP set in $G$. An EOP set with cardinality equal to $\rho^{o}_{e}(G)$ is called a $\rho^{o}_{e}$-set. It is noteworthy that if $D$ is an EOP set (resp. an induced matching), the induced subgraph $G[D]$ is a disjoint union of induced stars (resp. $K_{1,1}$-stars). Thus, the induced matching problem can be viewed as a special case of the EOP problem, which relaxes the stricter conditions of a matching or induced matching.
@@
An injective $k$-edge coloring of a graph $G$ is an assignment of colors (integers) from the set $\{1,\ldots,k\}$ to the edges of $G$ such that any two edges $e_{1},e_{2}\in E(G)$ that share a common adjacent edge, meaning an edge $e$ exists with one endpoint common to $e_{1}$ and the other endpoint common to $e_{2}$, are assigned distinct colors. The smallest integer $k$ for which this coloring exists is known as the injective chromatic index of $G$, denoted by $\chi^{\prime}_{i}(G)$. This notion was introduced by Cardoso et al. [cardoso2019injective] and subsequently explored in greater depth in [foucaud2021complexity, miao2022note]. It is noteworthy that each color class generated by an injective edge coloring constitutes an edge open packing (EOP) set. Consequently, determining $\chi^{\prime}_{i}(G)$ is effectively the problem of partitioning the edge set $E(G)$ into EOP sets. Practical applications of injective edge colorings and edge open packings have been discussed in the literature, specifically in [cardoso2019injective] and [chelladurai2022edge].
@@
In [chelladurai2022edge], Chelladurai et al. studied edge open packing in graphs. They showed some bounds in terms of a graph’s diameter, size, minimum degree, clique number, and girth. Also, they characterized graphs $G$ with $\rho_{e}^{o}(G)\in\{1,2,m-2,m-1,m\}$, where $m$ is the number of edges of $G$. Further, Brešar and Samadi [brevsar2024edge] study the EOP set problem from an algorithmic point of view. They proved that the decision problem associated with Maximum Edge Open Packing Problem is NP-complete for three special families of graphs, namely, graphs with universal vertices, Eulerian bipartite graphs, and planar graphs with maximum degree $4$. Moreover, they designed a linear-time algorithm in trees to solve the Maximum Edge Open Packing Problem. Also, they provided a structural characterization of all graphs $G$ attaining the upper bound $\rho_{e}^{o}(G)\leq|E(G)|/\delta(G)$. They analyzed the effect of edge removal on the EOP number of a graph $G$, obtaining a lower and an upper bound for $\rho_{e}^{o}(G-e)$, where $e$ is an edge of $G$. In [BRESAR2025], Brešar et al. showed some relationship between the induced matching number and the EOP number in trees and graph products. Later, Pandey and Santra [pandey2025edge] further characterize this parameter for any $t\geq 3$. They showed the necessary and sufficient condition for $\rho_{e}^{o}=t$, for $t\geq 3$, and explicitly defined graph families for $\rho_{e}^{o}=m-3$, where $m$ is the number of edges of $G$. In[brevsar2024edge], Brešar and Samadi pose the following open question of the edge open packing problem in chordal graphs.
@@
An edge between two non-consecutive vertices of a cycle is called a chord. If every cycle in $G$ of length at least four has a chord, then $G$ is called a chordal graph. A vertex $v\in V$ is called a simplicial vertex of $G$ if $N_{G}[v]$ induces a clique in $G$. A perfect elimination ordering (PEO) of $G$ is an ordering of the vertices, say $\sigma=(v_{1},v_{2},\ldots,v_{n})$, such that $v_{i}$ is a simplicial vertex of $G_{i}=G[\{v_{i},v_{i+1},\ldots,v_{n}\}]$ for all $1\leq i\leq n$. The existence of PEO can characterize chordal graphs; that is, a graph $G$ is chordal if and only if $G$ has a PEO [fulkerson1965incidence].
@@
For a chordal graph $G$, a perfect elimination ordering (PEO) $\alpha=(v_{1},v_{2},\ldots,v_{n})$ is called a bi-compatible elimination ordering (BCO) if the reverse ordering $\alpha^{-1}=(v_{n},v_{n-1},\ldots,v_{1})$ is also a PEO of $G$. A graph $G$ admits a BCO if and only if it is a proper interval graph [jamison1982elimination]. The following observation concerns the structure of BCOs in proper interval graphs.
Observation 2 ([panda2003linear]).
@@ 2 EOP set in proper interval graphs
Input: A connected proper interval graph $G=(V,E)$ with $n$ vertices.
@@
2Compute a BCO $\sigma=(v_{1},v_{2},\dots,v_{n})$ of $G$;
3 Let $k\leftarrow l(v_{1})$;
4
5for $j\leftarrow 2$ to $k$ do
6   let $\mathcal{H}_{j}\leftarrow G[v_{\,l(v_{j})+1},\dots,v_{n}]$;
 $\rho_{e}^{o}(\mathcal{H}_{j})\leftarrow\,$MaxEOP_PIG($\mathcal{H}_{j}$);
 // recursive call
7 if $l(v_{j})\neq k$ then
8    for $p\leftarrow k+1$ to $l(v_{j})$ do
9         let $\mathcal{G}^{p}_{j}\leftarrow G[v_{\,l(v_{p})+1},\dots,v_{n}]$;
       $\rho_{e}^{o}(\mathcal{G}^{p}_{j})\leftarrow\,$MaxEOP_PIG($\mathcal{G}^{p}_{j}$);
       // recursive call
@@ 3 Let $k\leftarrow l(v_{1})$;
[…]
15 $\rho^{\prime\prime}(G)\leftarrow 2+\displaystyle\max_{\begin{subarray}{c}2\leq j\leq k\\ k<p\leq l(v_{j})\end{subarray}}\rho_{e}^{o}(\mathcal{G}^{p}_{j})$;
16 $\rho_{1}(G)\leftarrow\max\{\rho^{\prime}(G),\rho^{\prime\prime}(G)\}$;
@@
18let $\mathcal{G}^{\prime}\leftarrow G[v_{2},v_{3},\dots,v_{n}]$;
19 $\rho_{2}(G)\leftarrow\rho_{e}^{o}(\mathcal{G}^{\prime})$;
20
[…]
@@ 19 $\rho_{2}(G)\leftarrow\rho_{e}^{o}(\mathcal{G}^{\prime})$;
23return $\rho_{e}^{o}(G)$;
@@ 3 EOP set in block graphs
In this section, we design a polynomial-time algorithm for finding the edge open packing number of a given block graph. We start with some pertinent definitions related to block graphs. Let $G$ be a block graph with $l$ blocks, $B_{1},B_{2},\ldots,B_{l}$, and $k$ cut vertices, $v_{1},v_{2},\ldots,v_{k}$. The cut-tree of $G$, denoted by $T_{G}=(V^{\prime},E^{\prime})$, is defined as $V^{\prime}=\{B_{1},B_{2},\ldots,B_{l},v_{1},v_{2},\ldots,v_{k}\}$ and $E^{\prime}=\{B_{i}v_{j}|v_{j}\in B_{i},1\leq i\leq l,1\leq j\leq k\}$. An example of a block graph and its cut-tree is illustrated in Figure 1. A vertex of $T_{G}$ corresponding to a block of $G$ is called a block vertex. The cut-tree $T_{G}$ of block graph $G$ can be made a rooted tree by selecting any block or cut vertex as the root. The set of all blocks of a given block graph $G=(V,E)$ can be found in $O(|V|+|E|)$ time by performing a depth-first search (DFS). For each cut vertex $v$ of $G$, the set of all blocks containing $v$ can also be found in $O(|V|+|E|)$ time. Hence the rooted cut-tree of a block graph can be constructed in $O(|V|+|E|)$ time [AhoHopcroftUllman].
@@ 5 Conclusion
In this paper, we have partially solved the open question posed by Brešar and Samadi about edge open packing sets in chordal graphs. First, we have designed a polynomial-time algorithm to solve the maximum edge open packing set in a proper interval graph. Furthermore, we have presented a polynomial-time algorithm to solve the maximum edge open packing set in block graphs. Finally, we have shown that the problem of edge open packing number can be solved in linear time in split graphs. It would be fascinating to investigate the other open problems mentioned in [chelladurai2022edge] and [brevsar2024edge].